# Exploring the Impact of Nuclear Parton Distribution Functions on Open Heavy Flavor Azimuthal Angular Correlations in Pb-Pb Collisions at $\sqrt{s_{NN}}$ = 5.5 TeV

### v1

@@ III.2 Methodology
In addition to the default proton PDF in PYTHIA8, we append LHAPDF6 to access the recent CT18Anlo Hou:2019efy proton PDF and the EPPS21nlo and nNNPDF3 nuclear PDFs. To establish the effect of nPDFs, we examine the production of Z-boson as a function of y. We deploy EPPS21, and nNNPDF3.0, which are modified free proton PDFs in the presence of a nuclear environment. In FIG 1, the ratio between these two PDF settings is compared with the Z-boson production data from the ATLAS ATLAS:2019maq and ALICE ALICE:2020jff experiments. To produce Z-bosons, weak processes were enabled using WeakSingleBoson:all=ON, WeakDoubleBoson:all=ON, and WeakBosonAndParton:all=ON. The y dependence also shows that at large rapidities nuclear PDFs provide a better description of experimental data than free proton PDF calculations.
Figure 1: The ratio of $Z$-boson yield (black points) as a function of rapidity measured by ALICE ALICE:2020jff and ATLAS ATLAS:2019maq in Pb-Pb collisions at $\sqrt{s}=5.02$ TeV to free-nucleon PDF (CT14). The ratios are compared to predictions using different free proton PDFs (CT14MC1nlo, CT18nlo) and nuclear PDFs (EPPS16, EPPS21).
@@
Figure 2: Comparison of azimuthal angular correlations of prompt $\rm{D^{0}-\bar{D}^{0}}$ mesons originating from the same hard scattering process, using free parton distribution functions (NNPDF2.3) and nuclear parton distribution functions (EPPS21, nNNPDF3), simulated with PYTHIA8+Angantyr at $\rm{\sqrt{s}=5.5\,TeV}$ is shown for different $p_{\rm{T}}^{\rm trig}$ and $p_{\rm{T}}^{\rm assoc}$ ranges.
Figure 3: Comparison of near-side yield (top row) and width (bottom row) of azimuthal angular correlation of $\rm{D^{0}-\bar{D}^{0}}$ meson using free parton distribution functions (NNPDF2.3) and nuclear parton distribution functions (EPPS21, nNNPDF3), simulated with PYTHIA8+Angantyr at $\rm{\sqrt{s}=5.5\,TeV}$. Triangular markers (hollow and filled) for $p_{\rm{T}}^{\rm trig}$ class I represent individual widths of two peaks, and square marker represents combined width.
Figure 4: Comparison of away-side yield (top row) and width (bottom row) of azimuthal angular correlation of $\rm{D^{0}-\bar{D}^{0}}$ meson using free parton distribution functions (NNPDF2.3) and nuclear parton distribution functions (EPPS21, nNNPDF3), simulated with PYTHIA8+Angantyr at $\rm{\sqrt{s}=5.5\,TeV}$.
@@
We generate $\sim$1B events for each PDF in Pb-Pb collisions at $\sqrt{s_{NN}}=$ 5.5 TeV. Heavy-flavor hadrons are selected by maintaining the ALICE3 acceptance of $\mathrm{|y|<4.0}$. The non-diffractive component of the total cross-section for all soft QCD processes is used with the switch SoftQCD: Non-diffractive=on with MPI, ISR, and FSR to simulate minimum bias events. In the case of prompt $\rm{D^{0}-\bar{D}^{0}}$, the correlation distribution was obtained by correlating $\mathrm{D^{0}}$ and $\mathrm{\bar{D}^{0}}$ coming from the same hard scattering. This is ensured by tracing the produced (anti-)charm quark until it hadronizes to $\mathrm{(\bar{D}^{0})D^{0}}$, through its decay daughter at each stage of fragmentation. To ensure they come from the same scattering, the mothers of $\mathrm{c}$ and $\mathrm{\bar{c}}$ are matched. The same procedure is repeated for $\mathrm{B-\bar{B}}$ meson correlation, but in this case, all B-meson species are used as trigger particles instead of just $\mathrm{B^{0}}$. In the case of non-prompt $\mathrm{D^{0},D^{+},D^{+}_{s}}$, and their antiparticles, we trace the $\mathrm{b}$ quark until it decays to one of the $\mathrm{D}$-meson species. Later, the mothers of $\mathrm{b}$ and $\mathrm{\bar{b}}$ are checked to ensure they come from the same scattering.
@@
$+\frac{Y_{\rm{AS}}}{\sqrt{2\pi}\sigma_{\rm{AS}}}\times e^{-\left(\frac{\Delta\phi-\pi}{\sqrt{2}\sigma_{\rm{AS}}}\right)^{2}}.$ (5)
Here, $Y_{\rm{NS}}$ and $Y_{\rm{AS}}$ are the yields of NS and AS peaks, $\beta_{\rm{NS}}$ is called shape parameter for the near-side peak, and $\alpha_{\rm{NS}}$ is related to the $\sigma_{\rm{NS}}$ (width) of the peak as
@@
The non-prompt D-meson correlation distribution is fitted with a single generalized Gaussian function for the NS peak, von Mises for the AS peak, and a $0^{\text{th}}$ order polynomial for the baseline identification given by
${f(\Delta\phi)}=b+\frac{Y_{\rm{NS}}\times\beta_{\rm{NS}}}{2\alpha_{\rm{NS}}\Gamma{(1/\beta_{\rm{NS}})}}\times e^{-\left(\frac{\Delta\phi}{\alpha_{\rm{NS}}}\right)^{\beta_{\rm{NS}}}}$
$+\frac{e^{\kappa_{AS}\cos(\Delta\phi-\pi)}}{2\pi I_{0}(\kappa_{AS})}$ (7)
where $b$ is the baseline, $\rm{\kappa}$ is the reciprocal measure of dispersion that quantifies the width of the distribution, $I_{0}$ is the $\rm{0^{\text{th}}}$ order modified Bessel function. The mean of the AS peak is fixed at $\rm{\Delta\phi=\pi}$. The NS width is obtained in a similar manner as prompt $\rm{D^{0}}$, but the AS width is estimated by calculating the sigma ($\mathrm{\sigma}$) from the von Mises function as per the given relation:
$\sigma=\sqrt{-2\log{\frac{I_{1}(\kappa)}{I_{0}(\kappa)}}}$ (8)
Here, $I_{0}$ and $I_{1}$ are the modified Bessel functions of $\rm{0^{\text{th}}}$ and $\rm{1^{\text{st}}}$ order, respectively, and $\kappa$ is measured by the von Mises function fit parameter. The error in the width ($\rm{d\sigma}$) is propagated by the relation:
$\text{d}\sigma=\frac{1}{\sigma}\left(\frac{I_{1}}{I_{0}}-\frac{I_{0}}{I_{1}}+\frac{1}{\kappa}\right)\text{d}\kappa$ (9)
where $\rm{\text{d}\kappa}$ is the uncertainty in $\rm{\kappa}$, obtained from the fit result.
In the case of $\rm{B-\bar{B}}$, the AS peak exhibits a Gaussian-like structure for low associated $p_{\rm{T}}$ ranges, while a triangular nature for high associated $p_{\rm{T}}$ ranges, which inspired us to develop a novel function for the away-side peak. The $\rm{B-\bar{B}}$ $\rm{\Delta\varphi}$ distribution is fitted with two generalized Gaussian functions for the NS peak, the convolution of a Gaussian function and a triangular function for the AS peak, and a $\mathrm{0^{\text{th}}}$ order polynomial for the baseline identification given by
@@
Here, the away-side mean is fixed at $\rm{\Delta\phi=\pi}$, and the parameter ``$a$” is responsible for the shape of the distribution. For small values of ``$a$”, the AS function behaves like a Gaussian, while for large values of ``$a$”, it becomes triangular. The width of the AS distribution is extracted from its standard deviation.
@@ IV Results
Figure 5: Comparison of azimuthal angular correlations of prompt $\rm{B-\bar{B}}$ mesons originating from the same hard scattering process, using free parton distribution functions (NNPDF2.3) and nuclear parton distribution functions (EPPS21, nNNPDF3), simulated with PYTHIA8+Angantyr at s= $\rm{\sqrt{s}=\,5.5TeV}$ is shown for different $p_{\rm{T}}^{\rm trig}$ and $p_{\rm{T}}^{\rm assoc}$ ranges.
Figure 6: Comparison of near-side azimuthal angular correlation of $\rm{B-\bar{B}}$ meson between free proton parton distribution function (NNPDF2.3) and nuclear PDF (EPPS21,nNNPDF3) with PYTHIA8+Angantyr in Pb-Pb collisions at $\rm{\sqrt{s}=5.5\,TeV}$ is shown for different $p_{\rm{T}}^{\rm trig}$ and $p_{\rm{T}}^{\rm assoc}$ ranges. The near-side peak is plotted for $\rm{|\Delta\phi|}$ = 1.
Figure 7: Comparison of near-side yield (top row) and sigma (bottom row) of azimuthal angular correlation of $\rm{B-\bar{B}}$ meson between free proton parton distribution function (NNPDF2.3) and nuclear PDF (EPPS21,nNNPDF3) with PYTHIA8+Angantyr in Pb-Pb collisions at $\rm{\sqrt{s}=5.5\,TeV}$ for different $p_{\rm{T}}^{\rm trig}$ and $p_{\rm{T}}^{\rm assoc}$ ranges.
Figure 8: Comparing away-side yield (top row) and sigma (bottom row) of azimuthal angular correlation of $\rm{B-\bar{B}}$ meson between free proton parton distribution function (NNPDF2.3) and nuclear PDF (EPPS21,nNNPDF3) with PYTHIA8+Angantyr in Pb-Pb collisions at $\rm{\sqrt{s}=5.5\,TeV}$ for different $p_{\rm{T}}^{\rm trig}$ and $p_{\rm{T}}^{\rm assoc}$ ranges.
Figure 9: Comparison of azimuthal angular correlation of non-prompt $\rm{D-\bar{D}}$ meson between free proton parton distribution function (NNPDF2.3) and nuclear PDF (EPPS21,nNNPDF3) with PYTHIA8+Angantyr in Pb-Pb collisions at $\rm{\sqrt{s}=5.5\,TeV}$.
Figure 10: Comparison of near-side yield (top row) and sigma (bottom row) for azimuthal angular correlation of non-prompt $\rm{D-\bar{D}}$ meson between free proton parton distribution function (NNPDF2.3) and nuclear PDF (EPPS21,nNNPDF3) with PYTHIA8+Angantyr in Pb-Pb collisions at $\rm{\sqrt{s}=5.5\,TeV}$.
Figure 11: Comparison of away-side yield (top row) and sigma (bottom row) for azimuthal angular correlation of non-prompt $\rm{D-\bar{D}}$ meson between free proton parton distribution function (NNPDF2.3) and nuclear PDF (EPPS21,nNNPDF3) with PYTHIA8+Angantyr in Pb-Pb collisions at $\rm{\sqrt{s}=5.5\,TeV}$.
Figure 12: Ratio of near-side yield (top row) and sigma (bottom row) of nuclear PDFs predictions to the free proton PDF predictions for $\rm{D^{0}-\bar{D}^{0}}$, $\rm{B-\bar{B}}$ and non-prompt $\rm{D-\bar{D}}$ azimuthal angular correlations. The left column shows the ratios for prompt $\rm{D^{0}}$ meson while the middle and right columns show the ratios for B-meson and non-prompt D-mesons respectively. Numbers on the x-axis represent different $p_{\rm{T}}^{\rm trig}$ classes while different shades of color represent various $p_{\rm{T}}^{\rm assoc}$ ranges.
Figure 13: Ratio of away-side yield (top row) and sigma (bottom row) of nuclear PDFs predictions to the free proton PDF predictions for $\rm{D^{0}-\bar{D}^{0}}$, $\rm{B-\bar{B}}$ and non-prompt $\rm{D-\bar{D}}$ azimuthal angular correlations. The left column shows the ratios for prompt $\rm{D^{0}}$ meson while the middle and right columns show the ratios for B-meson and non-prompt D-mesons respectively. Numbers on the x-axis represent different $p_{\rm{T}}^{\rm trig}$ classes, while different shades of color represent various $p_{\rm{T}}^{\rm assoc}$ ranges.
IV.1 $\rm{D^{0}-\bar{D}^{0}}$ azimuthal angular correlation
The azimuthal angular correlation of $\rm{D^{0}-\bar{D}^{0}}$ using NNPDF2.3, EPPS21, and nNNPDF3 for different triggers and associated $p_{\rm{T}}$ ranges is shown in FIG 2. The trigger particles are classified into 3 classes with $p_{\rm{T}}^{\rm trig}$ ranges $\mathrm{2-4}$, $\mathrm{\geq 4}$, and $\mathrm{\geq 6}$ $\mathrm{GeV/c}$, while associated particles are selected in the $p_{\rm{T}}^{\rm assoc}$ ranges of $\rm{2-4}$, $\rm{2-6}$, $\rm{2-50}\,GeV/c$. FIG 3 and FIG 4 provide a quantitative comparison of NS and AS yields and widths of the $\rm{\Delta\varphi}$ distribution, respectively. Key observations from prompt $\rm{D^{0}-\bar{D}^{0}}$ correlations are as follows:
  1. 1.
    The NS and AS peaks of the $\rm{D^{0}-\bar{D}^{0}}$ azimuthal angular correlation function for EPPS21 and nNNPDF3 are qualitatively similar in all $p_{\rm{T}}^{\rm trig}$ and $p_{\rm{T}}^{\rm assoc}$ intervals. In contrast, proton PDF correlations are distinctly different, showing a smaller yield and a double-peak structure at NS, which is unseen for nPDFs.
  2. 2.
    The angle between produced $\mathrm{c-\bar{c}}$ pairs during the gluon splitting process results in a double-peak structure at the NS. TABLE 1 shows that the separation between two NS peaks diminishes as the $p_{\rm{T}}^{\rm assoc}$ range increases for the given $p_{\rm{T}}^{\rm trig}$ range. The observation can be explained by the fact that, at low $p_{\rm{T}}^{\rm assoc}$, the opening angle is wide, while at higher $p_{\rm{T}}^{\rm assoc}$, the angle narrows due to a larger initial gluon boost. For $p_{\rm{T}}^{\rm trig}$ $>4$ GeV/$c$, the double NS peak disappears due to insignificant opening angles to resolve the two peaks.
    $p_{\rm{T}}^{\rm assoc}$ (GeV/$c$) Mean of near side peaks (rad)

### v2

@@ III.2 Methodology
In addition to the default proton PDF in PYTHIA8, we append LHAPDF6 to access the recent free proton PDFs (CT18ANLO Hou:2019efy , NNPDF4 NNPDF:2021njg ) and nuclear PDFs (EPPS21 Eskola:2021nhw , nNNPDF3 AbdulKhalek:2022fyi ). To establish the effect of nPDFs, we examine the production of Z-boson as a function of rapidity. In FIG 1, the ratio between these two PDF settings is compared with the Z-boson production data from the ATLAS and ALICE experiments. To produce Z-bosons, weak processes were enabled using WeakSingleBoson:all=ON, WeakDoubleBoson:all=ON. The rapidity dependence also shows that at large rapidities, nuclear PDFs provide a better description of experimental data than free proton PDF calculations.
@@
Figure 3: Azimuthal angular correlations of prompt $\rm{B-\bar{B}}$ mesons from the same hard scattering process in Pb-Pb collisions at $\sqrt{s}=5.5$ TeV, simulated using PYTHIA8+Angantyr with different PDF and nPDF settings. Results are shown for various $p_{\rm{T}}^{\rm B}$ and $p_{\rm{T}}^{\rm\bar{B}}$ ranges.
Figure 4: Comparison of near-side yield (top row) and width (bottom row) of azimuthal angular correlations for prompt $\rm{D^{0}-\bar{D}^{0}}$ mesons in Pb-Pb collisions at $\sqrt{s}=5.5$ TeV, simulated using PYTHIA8+Angantyr. Results are presented for different trigger momentum classes ($p_{\rm{T}}^{\rm D^{0}}$) and associated momentum ranges ($p_{\rm{T}}^{\rm\bar{D^{0}}}$), comparing various PDF and nPDF settings.
Figure 5: Comparison of away-side yield (top row) and width (bottom row) of azimuthal angular correlations for prompt $\rm{D^{0}-\bar{D}^{0}}$ mesons in Pb-Pb collisions at $\sqrt{s}=5.5$ TeV, simulated using PYTHIA8+Angantyr. Results are presented for different trigger momentum classes ($p_{\rm{T}}^{\rm D^{0}}$) and associated momentum ranges ($p_{\rm{T}}^{\rm\bar{D^{0}}}$), comparing various PDF and nPDF settings.
@@
We simulate Pb-Pb collisions at $\sqrt{s_{NN}}=$ 5.5 TeV for 6 combinations of PDFs. These combinations are NNPDF2.3, EPPS21 (Default MPI), EPPS21, EPPS21+NNPDF2.3, EPPS21+CT18ANLO, and nNNPDF3+NNPDF4. In the first three cases, the mentioned PDFs are used for hard scatterings, ISR, FSR, and MPIs, while in later cases, nPDFs are used for hard scatterings, and free proton PDFs are used for other processes. It has been observed that the choice of PDFs significantly changes the multiplicity distribution of an event. In all cases except for EPPS21 (Default MPI), multiplicity distribution is matched to ALICE midrapidity Pb-Pb data by tuning parameter related to MPIs in PYTHIA8 Bierlich:2022pfr . The non-diffractive component of the total cross-section for all soft QCD processes is used with the switch SoftQCD: Non-diffractive=on with MPI, ISR, and FSR to simulate minimum bias events. Open heavy-flavor trigger hadrons are selected by maintaining the ALICE3 acceptance of $\mathrm{|y|<4.0}$. In the case of prompt $\rm{D^{0}-\bar{D}^{0}}$, the correlation distribution was obtained by correlating $\mathrm{D^{0}}$ and $\mathrm{\bar{D}^{0}}$ coming from the same hard scattering. This is ensured by tracing the produced (anti-)charm quark until it hadronizes to $\mathrm{(\bar{D}^{0})D^{0}}$, through its decay daughter at each stage of fragmentation. To ensure they come from the same scattering, the mothers of $\mathrm{c}$ and $\mathrm{\bar{c}}$ are matched. The same procedure is repeated for $\mathrm{B-\bar{B}}$ meson correlation, but in this case, all B-meson species are used as trigger particles instead of just $\mathrm{B^{0}}$.
@@
$+\frac{Y_{\rm{AS}}}{\sqrt{2\pi}\sigma_{\rm{AS}}}\times e^{-\left(\frac{\Delta\phi-\pi}{\sqrt{2}\sigma_{\rm{AS}}}\right)^{2}}$ (5)
Here, $Y_{\rm{NS}}$ and $Y_{\rm{AS}}$ are the yields for NS and AS peaks, $\beta_{\rm{NS}}$ is the shape parameter for the NS peak, and $\alpha_{\rm{NS}}$ is related to the $\sigma_{\rm{NS}}$ (width) of the peak as
@@
In the case of $\rm{B-\bar{B}}$, the AS peak exhibits a Gaussian-like structure for low associated $p_{\rm{T}}$ ranges, while a triangular nature for high associated $p_{\rm{T}}$ ranges inspired us to develop a novel function for the AS peak. The $\rm{B-\bar{B}}$ $\rm{\Delta\varphi}$ distribution is fitted with two generalized Gaussian functions for the NS peak, the convolution of a Gaussian function and a triangular function for the AS peak, and a $\mathrm{0^{\text{th}}}$ order polynomial for the baseline identification given by
@@
Here, the AS mean is fixed at $\rm{\Delta\phi=\pi}$, and the parameter ``$a$” is responsible for the shape of the distribution. For small values of ``$a$”, the AS function behaves like a Gaussian, while for large values of ``$a$”, it becomes triangular. The width of the AS distribution is extracted from its standard deviation.
@@
Figure 6: Comparison of near-side yield (top row) and width (bottom row) of azimuthal angular correlations for prompt $\rm{D^{0}-\bar{D}^{0}}$ mesons in Pb-Pb collisions at $\sqrt{s}=5.5$ TeV, simulated using PYTHIA8+Angantyr. Results are presented for different trigger momentum classes ($p_{\rm{T}}^{\rm B}$) and associated momentum ranges ($p_{\rm{T}}^{\rm\bar{B}}$), comparing various PDF and nPDF settings.
Figure 7: Comparison of away-side yield (top row) and width (bottom row) of azimuthal angular correlations for prompt $\rm{D^{0}-\bar{D}^{0}}$ mesons in Pb-Pb collisions at $\sqrt{s}=5.5$ TeV, simulated using PYTHIA8+Angantyr. Results are presented for different trigger momentum classes ($p_{\rm{T}}^{\rm B}$) and associated momentum ranges ($p_{\rm{T}}^{\rm\bar{B}}$), comparing various PDF and nPDF settings.
Figure 8: Ratios of near-side yield (top) and width (bottom) of azimuthal angular correlations to the NNPDF2.3 baseline for prompt $\rm{D^{0}-\bar{D}^{0}}$, and $\rm{B-\bar{B}}$ meson pairs in Pb–Pb collisions at $\sqrt{s_{NN}}=5.5$ TeV, simulated using PYTHIA8+Angantyr for various PDF and nPDF settings. The left and right columns correspond to prompt $\rm{D^{0}}$, and B-mesons, respectively. Different line styles represent $p_{\rm{T}}^{\rm assoc}$ ranges, scaled for better visualization.
Figure 9: Ratios of away-side yield (top) and width (bottom) of azimuthal angular correlations to the NNPDF2.3 baseline for prompt $\rm{D^{0}-\bar{D}^{0}}$, and $\rm{B-\bar{B}}$ meson pairs in Pb–Pb collisions at $\sqrt{s_{NN}}=5.5$ TeV, simulated using PYTHIA8+Angantyr for various PDF and nPDF settings. The left and right columns correspond to prompt $\rm{D^{0}}$ and B-mesons, respectively. Different line styles represent $p_{\rm{T}}^{\rm assoc}$ ranges, scaled for better visualization.
@@ IV Results
The azimuthal angular correlation of $\rm{D^{0}-\bar{D}^{0}}$ pairs for different trigger and associated transverse momentum ($p_{\rm{T}}$) ranges is presented in FIG. 2. Based on the $p_{\rm{T}}$, $\rm{D^{0}}$’s are categorized into three classes: Class I ($p_{\rm{T}}^{\rm trig}$ = 2–4 GeV/c), Class II ($p_{\rm{T}}^{\rm trig}$geq 4 GeV/c), and Class III ($p_{\rm{T}}^{\rm trig}$geq 6 GeV/c). The associated particles are selected within the $p_{\rm{T}}^{\rm assoc}$ ranges of 2–4, 2–6, and 2–50 GeV/c. For B-mesons (FIG. 3), the AS distribution is found to dominate over the NS distribution, exhibiting a broader width at the lower trigger and associated $p_{\rm{T}}$ ranges. Therefore, to investigate the characteristics of azimuthal angular correlations of B mesons arising from different production processes, the trigger particle classes are defined with $p_{\rm{T}}^{\rm trig}$ ranges of 10–16, $\geq$ 16, and $\geq$ 24 GeV/c and the associated particles are selected within $|\Delta\eta|<1$ and $p_{\rm{T}}^{\rm assoc}$ ranges of 5–10, 10–15, and $>5$ GeV/c which are higher compared to $\rm{D^{0}-\bar{D}^{0}}$.
FIG. 2 and 3 show that the NS distribution is smaller than the AS distribution. This is because, at LHC energies, the cross-section for the gluon splitting process, which contributes to the NS distribution, is smaller compared to flavor excitation and pair creation, which predominantly contribute to the AS distribution.
$p_{\rm{T}}^{\rm assoc}$ (GeV/$c$) Mean of NS peaks (rad)